OpenStudy (anonymous):

A carnival game had a spinning wheel with 20 equal-size sections: some red, some black, and some green. Every time the wheel was spun, the game attendant recorded the color selected. The table shows the results after 20, 200, and 2000 spins. What is the best estimate of the number of green sections on the wheel? Outcome Number of trials 20 200 2000 red 12 109 1028 black 8 84 879 green 0 7 95

OpenStudy (anonymous):

A. 1 B. 2 C. 5 D. 9

OpenStudy (anonymous):

i am thinking either a or b

OpenStudy (misty1212):

HI!

OpenStudy (anonymous):

i will give medal and fan and perhaps testimony

OpenStudy (anonymous):

HEllo

OpenStudy (anonymous):

do u mind helping?

OpenStudy (anonymous):

helping me

OpenStudy (anonymous):

i am having a bit of trouble with this question

OpenStudy (misty1212):

i would try to solve \[\frac{95}{2000}=\frac{x}{20}\]

OpenStudy (anonymous):

i am not the best at math :P

OpenStudy (anonymous):

okay, i will tell u what i get in a sec...

OpenStudy (anonymous):

for solving x, i got 19

OpenStudy (misty1212):

solve that by \[95\times 20=2000x\] or \[x=\frac{95\times 20}{2000}\]

OpenStudy (misty1212):

you are off a bit. it should be \[\frac{19}{20}\]

OpenStudy (misty1212):

which is pretty close to 1

OpenStudy (anonymous):

oh, so it is a

OpenStudy (anonymous):

okay, oh i get it

OpenStudy (anonymous):

so i was right :)

OpenStudy (anonymous):

u in k12?

OpenStudy (anonymous):

What does it matter if she is in K12 or not? she's asking for help.

OpenStudy (anonymous):

calm down GOD